how do you solve(5a2/6b)(4b/a3)
Found 2 solutions by nyc_function, palanisamy:
Answer by nyc_function(2741)

Apply the same rules for regular fraction multiplication.
(5a^2/6b)(4b/a^3) = 10/3a

Answer by palanisamy(496)

(5a^2/6b)*(4b/a^3)=(20a^2b)/(6a^3b)=10/3a